OpenStudy (vane11):

I'm having issues understanding how to use the Arrhenius equation, usually on the step before the last where you do something with an exponent. Can someone help me understand?

OpenStudy (aaronq):

\(\Large k=A*e^{(\frac{-E_a}{RT})}\) divide both sides by the constant, A. \(\Large \dfrac{k}{A}=e^{(\frac{-E_a}{RT})}\) "take ln" of both sides \(\Large ln(\dfrac{k}{A})=ln(e^{(\frac{-E_a}{RT})})\) using logarithm rules, you can bring the exponent part down: \(\Large ln(\dfrac{k}{A})=(\frac{-E_a}{RT})ln(e)\) because \(ln = log_e\), \(ln(e)=log_e(e)=1\) you're left with: \(\Large ln(\dfrac{k}{A})=(\frac{-E_a}{RT})\) i think you can take it from here